# PET-MAD, a universal interatomic potential for advanced materials modeling

### v1

@@ Acknowledgements.
The Authors would like to thank Giovanni Pizzi, Marnik Bercx and Sebastian Huber for helping with the set up of AiiDA calculations on the MC3D dataset, and Sandip De for discussion on potentials for high-entropy alloys. They are also grateful to all the current and past members of the Laboratory of Computational Science and Modeling who contributed to the software infrastructure that supported this work. Computation for this work relied on resources from the Swiss National Supercomputing Centre (CSCS) under the projects s1092, s1219, s1243, s1287, Finnish IT Center for Science (CSC) under the project 465000551, and the EPFL HPC platform (SCITAS).
@@
AM and MC acknowledge support from an Industrial Grant from BASF and from EPFL. GF and MC acnowledge funding from the MARVEL National Centre of Competence in Research (NCCR), funded by the Swiss National Science Foundation (SNSF, grant number 182892) SP, FB and GF Platform were supported by a project within the Platform for Advanced Scientific Computing (PASC). DT acknowledges support from a Sinergia grant of the Swiss National Science Foundation (grant ID CRSII5_202296). PL, PP and MC acknowledge the funding from the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation programme (grant agreement No 101001890-FIAMMA). MK and MC acknowledge support by a Swiss National Science Foundation (grant ID 200020_214879).
@@
A.M. worked on creation of the MAD dataset, training the PET-MAD model, running the accuracy benchmarks for PET-MAD and reference models, performing the calculations of surface segregation in CoCrFeMnNi. F.B. worked on implementation of the metatrain infrastructure for training, evaluating the PET-MAD model, as well as performing the uncertainty quantification for it, performed the speed benchmarking for PET-MAD and reference models, and studied the nuclear quantum effects in water. M.K. calculated the melting point in GaAs and studied the chemical shielding in succinic acid. P.P. studied the dielectric response in barium titanate. D.T. studied the ionic conductivity in lithium triophosphate. G.F. developed the metatensor ecosystem, which metatrain infrastructure is based on, developed the interfaces for ASE and LAMMPS simulation engines, and helped with the computational setup for creating the MAD dataset. S.P. developed the original version of PET architecture. P.L. worked on implementation of the metatrain infrastructure and finalized the figures. M.C. designed and guided the project, and provided a theoretical support. All authors contributed to the original version of the manuscript.
@@
We compared the accuracy of PET-MAD against four recent universal machine learning interatomic potentials - MACE-MP-0 L, MatterSim-5M, Orb-v2, and SevenNet-l3i5 - on popular atomistic ML datasets: MPtrj, Alexandria, SPICE, MD22, and OC2020 S2EF. For each dataset (including MAD), we prepared a small subset of structures that were recalculated using DFT settings consistent with each model’s training data.
@@
The MAD benchmark consists of 322 structures selected by sampling 50 random structures from each MAD test subset, recalculated with MPtrj DFT settings, and cleaned of non-converged structures and outliers. The MPtrj benchmark is based on the MACE-MP-0 validation subset, reduced to 153 structures after removing four 1D wire structures. The OC2020-S2EF benchmark consists of 89 structures, where 100 structures were first sampled from the OC2020-S2EF training dataset and then cleaned of non-converged cases. The SPICE benchmark contains 100 randomly sampled neutral molecules from the SPICE dataset. The Alexandria benchmark includes 150 structures: 50 randomly sampled from Alexandria-2D and Alexandria-3D-gopt, and 100 from Alexandria-3D-relaxed. The MD22 benchmark consists of 134 structures, where 25 structures were first randomly sampled from each of the seven subsets of the original MD22 dataset (Ac-Ala3-NHMe, AT-AT, DHA, Stachyose, AT-AT-CG-CG, Buckyball-Catcher, double-walled-nanotube), and then cleaned of non-converged cases.
@@
S4 Diatomic energy curves
@@
$\sigma_{i}^{2}=\alpha\,\mathbf{f}_{i}^{\top}(\mathbf{F}^{\top}\mathbf{F}+\sigma^{2}\mathbf{I})^{-1}\mathbf{f}_{i}.$ (1)
In this expression, $\sigma_{i}$ is the uncertainty on the prediction relative to sample $i$, $\mathbf{f}_{i}$ are the latent features in the last layer of the neural network, $\mathbf{F}$ is a matrix whose rows correspond to each set of last-layer features for each structure in the training set, and $\sigma$ is a small regularizer. Besides computing energy uncertainties at nearly no additional cost compared to the raw predictions, the LLPR also allows the sampling of a last-layer ensemble [28], whose uncertainties can be propagated numerically through arbitrarily difficult workflows, including molecular dynamics [30]. The GaAs calculations in Section II.5 demonstrate how combining LLPR-sampled last-layer ensemble models with thermodynamic reweighting enables uncertainty quantification of PET-MAD predictions and propagate their uncertainties to thermodynamic averages at virtually no additional cost, at both training and inference time. The LLPR uncertainties, as well as those from an LLPR-derived last-layer ensemble of 128 members, are shown in Figure S4 for energy predictions on the MAD test set. The predictions follow the expected distribution almost exactly (see, for example, Refs. [28] and [30] for a discussion of this type of plots), confirming the very high quality of PET-MAD uncertainties, at least within the training domain.
S6 Fine-tuning accuracies
For each simulation case presented in this work we trained a bespoke PET model from scratch, and compared it against the LoRA-fine-tuned version. While being equally accurate in predicting observables, the fine-tuned model retains a certain degree of accuracy on the base MAD dataset, which can be beneficial in certain computational setups. In Table 2, we list the mean absolute errors of each fine-tuned model in predicting the energies and forces on the base MAD test set. For reference, the general-purpose PET-MAD model yields errors of 15.1 meV/at. | 72.3 meV/Å on the MAD test set.
Table 2: Accuracies of the LoRA-fine-tuned models in predicting the energies (in meV/at.) and forces (in meV/Å) on the MAD test set. For the pre-trained PET-MAD errors are 15.1 meV/at. | 72.3 meV/Å.
@@
S7.2 Ionic transport in lithium thiophosphate
Figure S6 shows learning curves for the Li3PS4 dataset for three learning scenarios, i.e., training from scratch (orange), LoRA fine-tuning with a rank of 8 (blue), and complete fine-tuning of all the model parameters (green). The black dashed line corresponds to the MSE of PET-MAD on the LPS test set. The complete fine-tuned seems to outperform the others of about 20 %, possibly because of the reduced number of epochs required during fine-tuning to reach a small error with respect to a model trained from scratch. LoRA fine-tuning provides slightly higher errors on forces with respect to the bespoke PET model, however from the results in the main text seems that this small difference does not translate in a large difference in the values of the ionic conductivity.
Figure S6: Learning curves (forces) for the Li3PS4 dataset of Ref. [42], comparing a bespoke model and a LoRA finetuned model with rank 8 against the general PET-MAD model.
S7.3 Melting point of GaAs
@@
S7.4 Surface segregation in high-entropy alloys
Figure S8 shows the learning curves of three models, namely the bespoke PET model trained on a subset of the HEA25S dataset [55] from scratch, and two fine-tuned models, which start from the pre-trained PET-MAD weights: the fully fine-tuned and a LoRA-fine-tuned model. Both fine-tuned models show similar training behavior in a low-data regime with a fully fune-tuned one demonstrating better accuracy as the amount of training data increases. In contrast to other simulations cases done in this work, the bespoke model cannot achieve the same accuracy on a test set, compared to fine-tuned models. This is likely caused by undersampling in the HEA25S subset used for training, which only contains 1975 structures with 25 transition metals, while the original dataset has about 30,000 structures.
@@
S7.5 Quantum nuclear effects in liquid water
Figure S9 shows learning curves for the liquid water dataset for three learning scenarios: (1) training from scratch, (2) LoRA fine-tuning with a rank of 8, (3) LoRA fine-tuning with a rank of 32. It can be seen that fine-tuning the PET-MAD model under these conditions is advantageous until approximately 20% of the entire training set (which consists of 1228 structures overall) is used. With more structures, the fine-tuned model is less accurate; this is partially remedied by increasing the rank of LoRA.
Figure S9: Learning curves for the water dataset, comparing training from scratch, LoRA fine-tuning with a rank of 8, and LoRA fine-tuning with a rank of 32 against the base PET-MAD model predictions error.
S7.6 Quantum nuclear effects in NMR crystallography
@@
S7.7 Dielectric response of barium titanate
Figure S12 shows learning curves for the BTO dataset for three learning scenarios, i.e., training from scratch, LoRA fine-tuning with a rank of 8, and fine-tuning of all the model parameters. Fine-tuned models always outperforms models trained from scratch, possibly because of the reduced number of epochs required during fine-tuning to reach a small error with respect to a model trained from scratch. LoRA fine-tuning provides slightly smaller errors on forces with respect to full fine-tuning, while at the same time being faster, requiring the update of a reduced number of parameters.
Figure S12: Learning curves for the BTO dataset, comparing training from scratch (PET-Bespoke) and LoRA fine-tuning with a rank of 8 (PET-MAD-LoRA) against the base PET-MAD model predictions error.
S8 Simulations
@@
S8.1 Ionic transport in lithium thiophosphate
The ionic conductivities, $\sigma$, in Section II.4 were computed via the Green-Kubo theory of linear response [44, 45], which is a practical framework to compute transport coefficients of extended systems [42]. For an isotropic system of $N$ interacting particles reads:
[…]
@@
Here, the sum runs over all the atoms, $e$ is the electron charge, and the $q_{i}$ are equal to the nominal oxidation number of the atoms [80].
S8.2 Melting point of GaAs
The melting points in Section II.5 were computed from interface pinning simulations of GaAs liquid solid interface model structures, following Imbalzano’s work, studying GaAs with MLIPs [47]. At the melting point, the chemical potential of the liquid phase $\mu_{l}$ and solid phase $\mu_{g}$ are identical. Applying a pinning potential restrains a system to coexisting liquid and solid phases separated by a planar interface. A suitable collective variable must be chosen that distinguishes between liquid and solid phase - we will employ the same atom centered Steinardt Q4 order parameter as chosen by Imbalzano. For a given configuration $A$, the expression for the pinning potential $V_{s}(A)$, reads:
@@
The melting point is then determined via root finding by simulating pinned systems at increasing temperatures and determining when the average force is zero. We compute the collective variable with PLUMED and run molecular dynamics with lammps to determine the average pinning force. Interface structures are constructed by first building elongated supercells (1152 Atoms, 17 Å$\times$ 17 Å$\times$ 90 Å) of GaAs and partial melting of half of the box, whilst constraining the solid half. These simulations were used to determine the average restraining value $\overline{s}$. The supercells were relaxed with the respective PET potential and then partially melted with NPT MD runs. Production MDs were run for 1 ns, using a 4 fs timestep. The Volume is kept fixed in the x and y directions while a barostat is applied on the z direction to fix the pressure at the interface around 1 bar. Simulations were performed for temperatures from 950 K to 1200 K in intervals of 50 K.
@@
S8.3 Surface segregation in high-entropy alloys
@@
S8.4 Quantum nuclear effects in liquid water
@@
The treatment of NQEs for equilibrium observables calculated with PIMD is exact as $P\rightarrow\infty$, and it reduces to that given by classical MD for $P=1$. The calculation of thermodynamic averages from PIMD simulations involves the use of so-called “estimators” [58], which can range from simple averages over the bead positions for position-dependent observables to more complicated expressions for momentum-dependent observables. In this work, we make use the heat capacity estimators from Ref. [59].
@@
S8.5 Quantum nuclear effects in NMR crystallography
@@
S8.6 Dielectric response of barium titanate

### v2

@@ Acknowledgements.
The Authors would like to thank Giovanni Pizzi, Marnik Bercx and Sebastian Huber for helping with the set-up of AiiDA calculations on the MC3D dataset, and Sandip De for discussion on potentials for high-entropy alloys. They are also grateful to all the current and past members of the Laboratory of Computational Science and Modeling who contributed to the software infrastructure that supported this work. Computation for this work relied on resources from the Swiss National Supercomputing Centre (CSCS) under the projects s1092, s1219, s1243, s1287, lp26, Finnish IT Center for Science (CSC) under the project 465000551, and the EPFL HPC platform (SCITAS).
@@
AM and MC acknowledge support from an Industrial Grant from BASF and from EPFL. GF and MC acknowledge funding from the MARVEL National Centre of Competence in Research (NCCR), funded by the Swiss National Science Foundation (SNSF, grant number 182892) SP, FB and GF Platform were supported by a project within the Platform for Advanced Scientific Computing (PASC). DT acknowledges support from a Sinergia grant of the Swiss National Science Foundation (grant ID CRSII5_202296). PL, PP and MC acknowledge the funding from the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation programme (grant agreement No 101001890-FIAMMA). MK and MC acknowledge support by a Swiss National Science Foundation (grant ID 200020_214879).
@@
A.M. worked on creation of the MAD dataset, training the PET-MAD model, running the accuracy benchmarks for PET-MAD and reference models, performing the calculations of surface segregation in CoCrFeMnNi. F.B. worked on implementation of the metatrain infrastructure for training and evaluating the PET-MAD model, as well as performing the uncertainty quantification for it, performed the speed benchmarking for PET-MAD and reference models, and studied the nuclear quantum effects in water. M.K. calculated the melting point in GaAs and studied the chemical shielding in succinic acid. P.P. studied the dielectric response in barium titanate. D.T. studied the ionic conductivity in lithium triophosphate. G.F. developed the metatensor ecosystem, which metatrain infrastructure is based on, developed the interfaces for ASE and LAMMPS simulation engines, and helped with the computational setup for creating the MAD dataset. S.P. developed the original version of PET architecture. P.L. worked on implementation of the metatrain infrastructure and finalized the figures. M.C. designed and guided the project, and provided a theoretical support. All authors contributed to the original version of the manuscript.
@@
We compared the accuracy of PET-MAD against four recent universal machine learning interatomic potentials - MACE-MP-0 L, MatterSim-5M, Orb-v2, and SevenNet-l3i5 - on popular atomistic ML datasets: MPtrj, Matbench Discovery, Alexandria, SPICE, MD22, and OC2020 S2EF. For each dataset (including MAD), we prepared a small subset of structures that were recalculated using DFT settings consistent with each model’s training data.
@@
The MAD benchmark consists of 360 structures selected by sampling 50 random structures from each MAD test subset, recalculated with MPtrj DFT settings, and cleaned of non-converged structures and outliers. The MPtrj benchmark is based on the MACE-MP-0 validation subset, reduced to 136 structures after removing four 1D wire structures. The Matbench Discovery benchmark contains 555 structures, randomly sampled from the original WBM dataset [16] - a part of the Matbench Discovery, which addresses the structural stability of inorganic crystals with random elemental substitutions. We didn’t include the structures with lanthanides and actinides to this subset in order to balance the effect of low coverage of these elements in the MAD dataset compared to other datasets like MPtrj. The OC2020-S2EF benchmark consists of 78 structures, where 100 structures were first sampled from the OC2020-S2EF training dataset and then cleaned of non-converged cases. The SPICE benchmark contains 99 structures, where 100 structures were randomly sampled neutral molecules from the SPICE dataset, and later cleaned of non-converges ones. The Alexandria benchmark includes 200 structures: 50 were randomly sampled from Alexandria-2D and Alexandria-3D-gopt, and 100 from Alexandria-3D. The MD22 benchmark consists of 134 structures, where 25 structures were first randomly sampled from each of the seven subsets of the original MD22 dataset (Ac-Ala3-NHMe, AT-AT, DHA, Stachyose, AT-AT-CG-CG, Buckyball-Catcher, double-walled-nanotube), and then cleaned of non-converged cases.
S3 Assessing PET-MAD on Matbench Discovery benchmark
Figure S2: Right panel: comparison of the energies above hull of a subset of structures from the WBM dataset [16], recomputed with DFT using MAD settings (MAD DFT, red dots), and predicted with PET-MAD (blue dots). The mean absolute error (MAE) in predicting the WBM values of the energy above hull in eV/atom is presented in the legend. This comparison demonstrates the upper limit of accuracy of any MAD-trained on the Matbench-Discovery benchmark upon using non-consistent DFT settings, as the difference in baseline DFT energy between Matbench and MAD settings reaches 120 meV/atom. Left panel: comparison of the PET-MAD predictions of the energy above hull against the consistent reference, recomputed with MAD settings. The resulting error decreases from 138 to 41 meV/atom, revealing the actual accuracy of the model.
As discussed throughout this work, it is crucial to maintain consistency in the level of DFT theory in the training set of the model. Lack of consistency in training data can introduce systematic errors to the model predictions, which can sometimes go up to 30 meV/atom (see Figure S3 for details). It is equally important to keep the consistency while assessing the accuracy on the model on benchmarks, like the popular Matbench Discovery [34]. Figure S2 demonstrates the results of the PET-MAD assessment on the WBM dataset [16] (a part of the Matbench Discovery) in predicting the energies above hull. When comparing the predictions against the default WBM values, PET-MAD yields a large error of around 140 meV/atom. However, this error is almost entirely explained by the significant discrepancy of 120 meV/atom in the underlying DFT reference, as demonstrated by comparing the WBM values to those recomputed using the MAD dataset settings on a subset of the WBM crystals. This value essentially sets an upper limit on the accuracy of any MAD-trained model when compared to data obtained with inconsistent (in this case - WBM) DFT settings. However, while using the consistent level of theory, one can significantly improve the models’ accuracy: in the left panel of the Fig. S2 we demonstrate the actual accuracy of the PET-MAD model if compared to the consistent DFT reference: the MAE drops from 138 meV/atom to 41 eV/atom.
This final MAE value was used to address another important property of the universal MLIPs - their data-efficiency. For all the models in Fig. 1 in the main text, we calculated the MAE value in predicting the energy above hull based on the same WBM subset, while utilizing the default WBM energy values for all the models except PET-MAD.
Therefore, since there is no straightforward way to account for significant discrepancies in the baseline DFT data in the Matbench Discovery benchmark, we decided not to add PET-MAD to the models list on the benchmark website. The resulting accuracy numbers would not be representative due to the lack of consistent DFT reference. We also didn’t include the phonon properties related part of the Matbench Discovery benchmark in this discussion, as we instead performed a more systematic study of the phonon properties in Supplementary Section S8.
S4 Effect of ab initio calculations settings convergence
@@
S5 Diatomic energy curves
@@
S6 Geometry optimization with PET-MAD
Figure S5: Timing distributions for 1000 geometry optimization runs using the L-BFGS optimizer. The black dashed band represents PET-MAD timings within three $\sigma$ from the median under the Gaussian approximation (0.135th to 99.865th percentile). The orange circles are the number of times the optimization runs did not complete.
We evaluated PET-MAD for geometry optimization on the three materials from section S8, namely, wurtzite BeO, zincblende BeTe, and rocksalt LiBr taken from the phononDB dataset. We used a simple optimization workflow using ASE calculators and the LBFGS optimizer with a default maximum force threshold of 0.05 eV/Å. To collect statistics, we performed 1000 optimization tasks per material, applying random displacements to atomic positions and lattice vectors from a normal distribution with zero mean and standard deviation of 0.1 Å. We measured the time to complete each optimization for the five universal MLIPs used in our analyses and show the time distributions in Figure S5. We limited the maximum number of optimization steps to 100. PET-MAD timings are comparable to or faster than those of all conservative models tested. The non-conservative Orb-v2 is consistently faster but less stable, being the only model that failed to converge in up to 10% of the displacements.
S7 Uncertainties of PET-MAD
@@
$\sigma_{i}^{2}=\alpha\,\mathbf{f}_{i}^{\top}(\mathbf{F}^{\top}\mathbf{F}+\varepsilon^{2}\mathbf{I})^{-1}\mathbf{f}_{i}.$ (1)
In this expression, $\sigma_{i}$ is the uncertainty on the prediction relative to sample $i$, $\mathbf{f}_{i}$ are the latent features in the last layer of the neural network, $\mathbf{F}$ is a matrix whose rows correspond to each set of last-layer features for each structure in the training set, and $\varepsilon$ is a small regularizer. Besides computing energy uncertainties at nearly no additional cost compared to the raw predictions, the LLPR also allows the sampling of a last-layer ensemble [30], whose uncertainties can be propagated numerically through arbitrarily difficult workflows, including molecular dynamics [32]. The GaAs calculations in Section II.5 demonstrate how combining LLPR-sampled last-layer ensemble models with thermodynamic reweighting enables uncertainty quantification of PET-MAD predictions and propagate their uncertainties to thermodynamic averages at virtually no additional cost, at both training and inference time. The LLPR uncertainties, as well as those from an LLPR-derived last-layer ensemble of 128 members, are shown in Figure S6 for energy predictions on the MAD test set. The predictions follow the expected distribution almost exactly (see, for example, Refs. [30] and [32] for a discussion of this type of plots), confirming the very high quality of PET-MAD uncertainties, at least within the training domain.
S8 Uncertainty estimation for phonon dispersion curves
Figure S7: LLPR ensembles of phonon bands (orange, thin lines) for three representative structures of the phononDB dataset, compared with reference DFT results (black, dashed lines). Average values of the bands are reported denoted by orange, thick lines.
In addition to complex properties that require MD sampling, such as the melting temperature discussed in the main text, uncertainty can also be estimated for static properties like phonon band structures. In Fig. S7, we present the LLPR ensemble of phonon bands for three representative materials from the PBEsol phononDB dataset [70]: wurtzite BeO, zincblende BeTe, and rocksalt LiBr, which were also analyzed in Ref. 71 using the MACE-MP-0 foundation model.
In the case of BeO, the optical bands appear slightly softer than in the reference calculations, which we checked to produce phonon bands consistent with calculations done with MAD settings with a frequency root mean square deviation of 3 cm-1. This behavior, however, is commonly observed for universal MLIPs, as shown in Ref. [72], and occurs for BeO as well as for the other materials presented here when using MACE-MP-0 [71]. The deviation is reflected in the UQ, which shows increased variance in the phonon band ensemble for optical modes.
S9 Fine-tuning accuracies
For each simulation case presented in this work we trained a bespoke PET model from scratch, and compared it against the LoRA-finetuned version. While being equally accurate in predicting observables, the fine-tuned model retains a certain degree of accuracy on the base MAD dataset, which can be beneficial in certain computational setups. In Table 2, we list the mean absolute errors of each fine-tuned model in predicting the energies and forces on the base MAD test set. For reference, the general-purpose PET-MAD model yields errors of 15.1 meV/at. | 72.3 meV/Å on the MAD test set.
Table 2: Accuracies of the LoRA-finetuned models in predicting the energies (in meV/at.) and forces (in meV/Å) on the MAD test set. For the pre-trained PET-MAD errors are 15.1 meV/at. | 72.3 meV/Å.
@@
S10.2 Ionic transport in lithium thiophosphate
Figure S9 shows learning curves for the Li3PS4 dataset for three learning scenarios, i.e., training from scratch (orange), LoRA finetuning with a rank of 8 (blue), and complete fine-tuning of all the model parameters (green). The black dashed line corresponds to the MSE of PET-MAD on the LPS test set. The complete fine-tuned seems to outperform the others of about 20 %, possibly because of the reduced number of epochs required during fine-tuning to reach a small error with respect to a model trained from scratch. LoRA finetuning provides slightly higher errors on forces with respect to the bespoke PET model, however from the results in the main text seems that this small difference does not translate in a large difference in the values of the ionic conductivity.
Figure S9: Learning curves (forces) for the Li3PS4 dataset of Ref. [43], comparing a bespoke model and a LoRA-finetuned model with rank 8 against the general PET-MAD model.
S10.3 Melting point of GaAs
@@
S10.4 Surface segregation in high-entropy alloys
Figure S11 shows the learning curves of three models, namely the bespoke PET model trained on a subset of the HEA25S dataset [50] from scratch, and two fine-tuned models, which start from the pre-trained PET-MAD weights: the fully fine-tuned and a LoRA-finetuned model. Both fine-tuned models show similar training behavior in a low-data regime with a fully funetuned one demonstrating better accuracy as the amount of training data increases. In contrast to other simulations cases done in this work, the bespoke model cannot achieve the same accuracy on a test set, compared to fine-tuned models. This is likely caused by undersampling in the HEA25S subset used for training, which only contains 1975 structures with 25 transition metals, while the original dataset has about 30,000 structures.
@@
S10.5 Quantum nuclear effects in liquid water
Figure S12 shows learning curves for the liquid water dataset for three learning scenarios: (1) training from scratch, (2) LoRA finetuning with a rank of 8, (3) LoRA finetuning with a rank of 32. It can be seen that fine-tuning the PET-MAD model under these conditions is advantageous until approximately 20% of the entire training set (which consists of 1228 structures overall) is used. With more structures, the fine-tuned model is less accurate; this is partially remedied by increasing the rank of LoRA.
Figure S12: Learning curves for the water dataset, comparing training from scratch, LoRA finetuning with a rank of 8, and LoRA finetuning with a rank of 32 against the base PET-MAD model predictions error.
S10.6 Quantum nuclear effects in NMR crystallography
@@
S10.7 Dielectric response of barium titanate
Figure S15 shows learning curves for the BTO dataset for three learning scenarios, i.e., training from scratch, LoRA finetuning with a rank of 8, and fine-tuning of all the model parameters. Fine-tuned models always outperforms models trained from scratch, possibly because of the reduced number of epochs required during fine-tuning to reach a small error with respect to a model trained from scratch. LoRA finetuning provides slightly smaller errors on forces with respect to full fine-tuning, while at the same time being faster, requiring the update of a reduced number of parameters.
Figure S15: Learning curves for the BTO dataset, comparing training from scratch (PET-Bespoke) and LoRA finetuning with a rank of 8 (PET-MAD-LoRA) against the base PET-MAD model predictions error.
S11 Simulations
@@
S11.1 Ionic transport in lithium thiophosphate
The ionic conductivities, $\sigma$, in Section II.4 were computed via the Green-Kubo theory of linear response [45], which is a practical framework to compute transport coefficients of extended systems [43]. For an isotropic system of $N$ interacting particles reads:
$\sigma=\frac{\Omega}{3k_{\mathrm{B}}T}\int_{0}^{\infty}\langle\mathbf{J}_{q}(\Gamma_{t})\cdot\mathbf{J}_{q}(\Gamma_{0})\rangle\,dt,$ (2)
where $k_{\mathrm{B}}$ is the Boltzmann constant, $T$ the temperature and $\Gamma_{t}$ indicates the time evolution of a point in phase space from the initial condition $\Gamma_{0}$, over which the average $\langle\cdot\rangle$ is performed. $\mathbf{J}_{q}$ is the charge flux, that depends only on the velocities of the atoms, $\mathbf{v}_{i}$, and their charges, $q_{i}$:
@@
Here, the sum runs over all the atoms, $e$ is the electron charge, and the $q_{i}$ are equal to the nominal oxidation number of the atoms.
S11.2 Melting point of GaAs
The melting points in Section II.5 were computed from interface pinning simulations of GaAs liquid solid interface model structures, following Imbalzano’s work, studying GaAs with MLIPs [47]. At the melting point, the chemical potential of the liquid phase $\mu_{l}$ and solid phase $\mu_{g}$ are identical. Applying a pinning potential restrains a system to coexisting liquid and solid phases separated by a planar interface. A suitable collective variable must be chosen that distinguishes between liquid and solid phase - we will employ the same atom centered Steinhardt Q4 order parameter as chosen by Imbalzano. For a given configuration $A$, the expression for the pinning potential $V_{s}(A)$, reads:
@@
The melting point is then determined via root finding by simulating pinned systems at increasing temperatures and determining when the average force is zero. We compute the collective variable with PLUMED and run molecular dynamics with lammps to determine the average pinning force. Interface structures are constructed by first building elongated supercells (1152 Atoms, 17 $\times$ 17 $\times$ 90 Å) of GaAs and partial melting of half of the box, whilst constraining the solid half. These simulations were used to determine the average restraining value $\overline{s}$. The supercells were relaxed with the respective PET potential and then partially melted with NPT MD runs. Production MDs were run for 1 ns, using a 4 fs timestep. The Volume is kept fixed in the x and y directions while a barostat is applied on the z direction to fix the pressure at the interface around 1 bar. Simulations were performed for temperatures from 950 K to 1200 K in intervals of 50 K.
@@
S11.3 Surface segregation in high-entropy alloys
@@
S11.4 Quantum nuclear effects in liquid water
@@
The treatment of NQEs for equilibrium observables calculated with PIMD is exact as $P\rightarrow\infty$, and it reduces to that given by classical MD for $P=1$. The calculation of thermodynamic averages from PIMD simulations involves the use of so-called “estimators” [55], which can range from simple averages over the bead positions for position-dependent observables to more complicated expressions for momentum-dependent observables. In this work, we use the heat capacity estimators from Ref. [56].
@@
S11.5 Quantum nuclear effects in NMR crystallography
@@
S11.6 Dielectric response of barium titanate
@@
S12 Non-conservative MD
We perform molecular-dynamics simulations of the molten phase of the ionic liquid 1-Butyl-3-methylimidazolium chloride (BMIM-Cl), starting from configurations equilibrated at $500$ K. We compare simulations using (i) the standard, conservative forces computed as derivatives of the PET-MAD energy head; (ii) the non-conservative, direct force prediction head (which is approximately 2 times faster); (iii) a multiple-time-stepping setup in which the conservative forces are evaluated every 8 steps to correct the non-conservative trajectory (which is approximately 1.8 times faster than conservative MD). We refer the reader to Ref. 33 for a general discussion of the advantages and problems of direct force prediction.
Figure S16: Kinetic temperature profiles along MD simulations, using conservative force models (blue), direct force heads (red) and multiple time stepping (green). (left) Constant-energy simulations initialized from a structure equilibrated at 500 K; note the exponential drift occurring with non-conservative forces. (middle) Constant-temperature simulations using stochastic velocity rescaling, with a thermostat relaxation time $\tau_{L}=$10 fs; direct-force simulations reach a steady-state with the overall temperature fluctuating around the target $T=500$ K. (right) Temperature profiles for individual atomic types (Cl and H shown) demonstrate that the direct-force simulations are problematic even with aggressive global thermostatting; Cl reach a steady-state temperature above 2000 K, which is compensated by H (and other) atoms being well below the average temperature. Note that the larger kinetic energy fluctuations for Cl atoms, seen also for conservative trajectories, is normal and consistent with the fact that only a few Cl ions are present in the simulation box.
Figure S17: Cl-Cl pair correlation function (left) and Cl mean square displacement (right) for simulations of BMIM-Cl at $T=500$ K, using conservative force models (blue), direct force heads (red) and multiple time stepping (green).
We first perform constant-energy simulations, using a time step of 0.5 fs and a velocity Verlet integrator. When using a conservative force calculation, the kinetic temperature of the atoms as a function of time (Fig. S16a) fluctuates around the initial equilibrium temperature, as expected. The non-conservative head leads to fast, catastrophic drift of the kinetic energy, that quickly leads to molecular dissociation and completely unphysical outcomes. A multiple-time-stepping strategy, instead, yields a stable trajectory, that is entirely consistent with conservative MD. As shown for other non-conservative potentials in Ref. 33, using an aggressive global thermostat avoids a drift in the overall temperature, but breaks energy conservation, leading to different chemical species (or more broadly, different degrees of freedom) reaching a different steady-state kinetic energy (Fig. S16b). Even in this case, a multiple-time-stepping simulation avoids these artifacts, while retaining most of the computational advantage.
Multiple time step calculations also yield static and dynamic properties that are consistent with those of a conservative calculation, as shown in Fig. S17. Both the Cl-Cl correlation function (reporting on the local structure in the liquid) and the Cl diffusion (reporting on the ionic conductivity) are within the statistical uncertainty. On the other hand, direct-force trajectories lead to extremely fast Cl diffusion (consistent with the much higher steady-state temperature) and the pair correlation function show the association of Cl dimers, another clear failure of the non-conservative sampling.